OpenStudy (anonymous):

For any positive number b not equal to 1 and any number or variable n, evaluate the following expression. logb(b^n) = wit logb being the base

OpenStudy (anonymous):

n

OpenStudy (anonymous):

how did u get it?

OpenStudy (anonymous):

Assuming n is positive: logb(b^n) = nlogb(b) and in general: loga(b) = log(b)/log(b) so in this case we have: nlogb(b) = n (log(b)/log(b)) = n

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

correction: and in general: loga(b) = log(b)/log(b) so in this case we have: I meant and in general loga(b) = log(b)/log(a)